# Spectral Persistent Homology: Persistence Signals

### 28 Dec 2023 v1

@@ Abstract
In this paper, we present a novel family of descriptors for persistence diagrams, reconceptualizing them as signals in +$\mathbb{R}^{2}_{+}$. This marks a significant advancement in Topological Data Analysis. Our methodology transforms persistence diagrams into a finite-dimensional vector space through functionals of the discrete measures induced by these diagrams. While our focus is primarily on frequency-based transformations, we do not restrict our approach exclusively to this types of techniques. We term this family of transformations as Persistence Signals and prove stability for some members of this family against the 1-Kantorovitch-Rubinstein metric, ensuring its responsiveness to subtle data variations. Extensive comparative analysis reveals that our descriptor performs competitively with the current state-of-art from the topological data analysis literature, and often surpasses, the existing methods. This research not only introduces a groundbreaking perspective for data scientists but also establishes a foundation for future innovations in applying persistence diagrams in data analysis and machine learning.
Keywords Topological Data Analysis  $\cdot$ Persistent Homology  $\cdot$ Spectral Persistence  $\cdot$ Frequency Transformations
@@ 1 Introduction
Topological Data Analysis (TDA) has emerged as a transformative approach in data science, providing unique insights into the underlying structure of complex datasets through the capture of their topological features. The effectiveness of machine learning algorithms, particularly in pattern recognition and feature extraction, underscores the importance of understanding data geometry. TDA offers a more sophisticated exploration of this geometric landscape, leading to numerous successful applications across various fields. Notable examples include neuroscience (Anderson et al. (2018), Dabaghian et al. (2012)), where TDA has also been instrumental in understanding complex neural networks and brain connectivity patterns. In the field of environmental science, TDA has been applied for climate data analysis, as seen in Nunez et al. (2019). Additionally, TDA has made significant contributions to materials science, particularly in the study of the structure of materials, as exemplified by Donato et al. (2016). These examples underscore the broad applicability and potential of TDA in enhancing our understanding of complex data across diverse scientific domains.
@@
Our contribution in this study is multi-faceted and significant in the field of TDA. We introduce a novel family of descriptors, termed Persistence Signals (PS), which represent a paradigm shift in how persistence diagrams are interpreted and utilized. By reconceptualizing persistence diagrams as signals within the $\mathbb{R}^{2}$ space, we enable their transformation into a finite-dimensional vector space. This is achieved through the application of functionals to the discrete measures induced by these diagrams. Our methodology is not limited to frequency-based techniques, allowing for a broader scope of analysis and application. A key aspect of our contribution is the proof of stability of some member of the PS family against the 1-Kantorovitch-Rubinstein metric. This crucial property underscores the robustness of our approach, ensuring that the PS descriptors are sensitive and responsive to subtle variations in data.
@@
Our approach not only provides a novel method for vectorizing persistence diagrams but also bridges two crucial and burgeoning realms: TDA and signal processing theory. This fusion opens the doors to an entirely new perspective in data analysis, offering inspiring ways to unravel complex data patterns and extract meaningful insights. Our methodology stands as a testament to the synergy between these two fields, enabling a more intuitive and accessible approach to understanding and manipulating TDA structures. As such, we see this work as a pioneering step into a new domain of study that we call Spectral Persistence Homology. To the best of our knowledge, this is the first instance where tools from classical signal processing theory have been employed to achieve a nearly lossless information vectorization method. Our method requires minimal tuning and is designed to be user-friendly, making it accessible even to those with limited knowledge of TDA. This groundbreaking approach not only simplifies the application of TDA but also expands its potential to a wider audience, promising significant advancements in various fields that rely on complex data analysis.
@@
The remainder of this paper is organized as follows: in Section 2, we review related work that connects TDA and ML, focusing particularly on classical methods of embedding persistence diagrams using vectorization or kernel methods. Section 3 delves into the background of persistent homology and persistent measures, offering foundational knowledge crucial for grasping the ideas from this work. In Section 4, we present our methodology along with the stability theorems that underpin it, providing theoretical insights into our approach. Section 5 is dedicated to showcasing our experimental results. Here, we not only present our findings but also compare them with standard methods from the TDA literature, highlighting the advancements our method brings. Finally, Section 6 is devoted to summarizing our key findings and outlining ideas for further work, suggesting potential avenues for future research and development based upon the groundwork established in this paper. The implementation of Persistence Signals has been made available to the public as an open-source resource [^1] https://github.com/majkevh/spectral-master.
@@ 3.1 Persistent Homology
A standard approach for illustrating the insights obtained from persistent homology is the use of a persistence diagram. In our study, we adopt the notation for persistence diagrams as delineated in Monod (2022). The detailed definition of persistence diagrams is provided in Definition 3.1.
@@
A persistence diagram $D$ is a locally finite multiset of points in the half-plane $\Omega=\{(x,y)\in$ $\left.\mathbb{R}^{2}\mid x<y\right\}$ together with points on the diagonal $\partial\Omega=\left\{(x,x)\in\mathbb{R}^{2}\right\}$ counted with infinite multiplicity. Points in $\Omega$ are called off-diagonal points. We denote as persistence of a feature, the value $y-x$.
@@
The collection of all persistence diagrams that possess finite $p$-total persistence is represented by $\mathcal{D}_{p}$. As mentioned in Monod (2022), the metric space $(\mathcal{D}_{p},\mathrm{W}_{p,q})$ occupies a pivotal role in TDA. Particularly noteworthy is the case when $q=\infty$. In this scenario, the metric space $(\mathcal{D}_{p},\mathrm{W}_{p,\infty})$ constitutes a complete and separable metric space—essentially, a Polish space—for any $1\leq p<\infty$. This characteristic endows the space with significant analytical power, as it ensures that statistical and probabilistic concepts such as probability measures, expectations, and variances are well-defined within this context, as elucidated by Mileyko et al. (2011).
@@ 3.2 Persistence Measures
In the realm of persistence measures, the concept of the optimal partial transport distance plays a pivotal role. This distance has been formally introduced by Figalli and Gigli (2010) and is detailed in Definition 3.6. In our study, we use the same notation as in Monod (2022).
@@
Building on this foundational concept, we can then consider the metric space defined as $\left(\mathcal{M}^{p},\mathrm{OT}_{p}\right)$. Here, $\mathcal{M}^{p}$ is equipped with the optimal partial transport distance $\mathrm{OT}_{p}$, serving as its metric. As established in Divol and Lacombe (2019), the subset $\mathcal{D}^{p}$ within $\mathcal{M}^{p}$ maintains its closed subspace property and inherits the key characteristics of $\mathcal{M}^{p}$. This includes being a Polish space, applicable for all $1\leq p<\infty$ and $1\leq q\leq\infty$. Consequently, the space $(\mathcal{M}_{p},\mathrm{OT}_{p,q})$ emerges as a robust framework for statistical and probabilistic applications. Moreover, within this framework, $\mathrm{OT}_{p,q}$ coincide with $\mathrm{W}_{p,q}$ when restricted to $\mathcal{D}_{p}$. As also mentioned in Monod (2022), this alignment underscores the role of persistence measures and the optimal partial transport distance as comprehensive extensions to traditional persistence diagrams and the $p$-Wasserstein distance.
@@ 4 Methodology
In this section, we introduce the Spectral Persistent Homology framework for embedding persistence measures in vector spaces along with the stability results.
As an overview, the process of vectorizing a persistence diagram using Persistence Signals, depicted in Algorithm 1, involves two distinct steps: first, we compute the discretized persistence diagram (refer to Section 4.2), transforming the persistence diagram into a format suitable for further processing. Following this, a functional is applied to this discretized representation to achieve an embedding within a vector space. As detailed in Section 4.1, this work primarily focuses on functionals that define valid frequency and time-frequency distribution functions. Notably, these include the Fourier Transform for frequency distribution, and the Wavelet and Gabor Transforms for time-frequency distribution. These transforms will be introduced in the subsequent subsection. It’s important to note, however, that our framework is versatile and not exclusively limited to these transforms.
Algorithm 1 Persistence Signals vectorization
1:$f$: Transformation function with hyperparameters, Grid size: $\{N,M\}$, PDs $\triangleright$ Section 4.1
@@ 4.1 Functionals on Persistence Measures
In this paper, we concentrate on functionals that establish valid frequency and time-frequency distribution functions (hereinafter referred to, for simplicity, as frequency transforms). Both families of these transforms are mapped onto the complex hyperplane (frequency domain from now on). Consequently, $\mathcal{F}(\mathcal{M}^{p})$ can be characterized as a vector space. The choice of selecting functionals that map into the frequency domain is motivated by the fact that persistent diagrams summarize topological features across various length scales of a filtration parameter. This parameter can be seen as a timestamp, indicating when topological features emerge or vanish, therefore it seems natural to interpret them as signals.
In this work, we focus on three possible choices of $f$: two time-frequency distributions, namely the Gabor Transform and the Wavelet Transform, and one frequency transform, specifically the Fourier Transform (which can be regarded as a special case of the Gabor Transform).
@@
4.1.1 Grid Based Approach in Computational Persistence Theory
Considering the practical computational requirements of all transforms under consideration, our theory is formulated based on the computation of discretized persistence diagrams (for details, refer to Section 4.2). In this framework, measures are confined to allocating mass exclusively at points on a grid (for a definition of the grid measure subspace $\Gamma_{p}$, refer to Lemma 4.1). Consequently, to compute the embedding, we employ a discretized version of the transforms discussed in the previous section. For a comprehensive literature review on the Discrete Fourier Transform (DFT), Discrete Wavelet Transform (DWT), and Discrete Gabor Transform (DGT), we direct the reader to Oppenheim and Schafer (2009) and Søndergaard (2007). Given a persistence measure $\mu\in\Gamma_{p}$, the three functionals on the transforms depicted in Section 4.1 are now formulated as follows:
@@ 4.2 Discretized Persistence Diagrams
The concept of discretization of persistence diagrams naturally arises from the grid-based algorithms used in real-life computations of the transforms outlined in Section 4.1, and it forms a critical aspect of our methodology.
@@ 4.2.2 Mapping Step
By Proposition 4.2, we therefore obtain a persistence measure $\nu^{\star}$ on $\Gamma_{p}$. We refer to $\nu^{\star}$ as the measure induced by the so-called discretized persistence diagram.
@@ 4.3 Stability
In line with this, we have established stability results with respect to the 1-Wasserstein metric for two of the transforms we considered, namely the Fourier and Gabor Transforms. For the Wavelet Transform, there does not always exist a closed-form expression, which makes finding stability results very challenging. Extending these results to the Wavelet Transform remains a challenge for future work.
@@
Let $\mu=\sum_{\mathbf{x}\in\Theta_{\tau}(D)}\delta_{\mathbf{x}}$ and $\nu=\sum_{\mathbf{x}\in\Theta_{\tau}(D^{\prime})}\delta_{\mathbf{x}}$ be two measures on the discretized persistence diagrams of $D,D^{\prime}$, let $f$ be a functional, either Gabor or Fourier (refer to Section 4.1). Then, there exists a finite constant $c$ such that:
@@ 5 Results
In this section, we demonstrate how PS effectively preserve topological information, proving to be as competitive as state-of-the-art methods from the TDA litarature. Our experiments are conducted on two types of classification tasks: graph-based classification and classification of point clouds derived from dynamical systems. In our experiments, we utilized a Vanilla Random Forest classifier from scikit-learn. The choice of a random forest classifier was driven by two primary considerations: Firstly, it facilitates a direct comparison with the study conducted by Royer et al. (2021). Secondly, in line with our objective to showcase the overall effectiveness of our methods relative to existing literature, employing this type of classifier helps to circumvent the necessity for elaborate infrastructure or extensive fine-tuning. Such an approach ensures that the results are not excessively influenced by particular design choices.
@@ 5.1 Graph Classification
The highest accuracies achieved with these transforms are presented in Table 1. Detailed observations for each transform across the different grids can be found in the Appendix B.Performance is assessed through 10-fold evaluations on each dataset, reporting both average and best 10-fold results.
@@
Dataset ${}^{\dagger}$ ${}^{\star}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\star}$ ${}^{\star}$ ${}^{\dagger}$ PS
${}^{\star}$ ${}^{\dagger}$
MUTAG $88.3$ $90.3$ $92.1$ $86.7$ $89.0$ $79.2$ $86.1$ $89.8$ $88.3$ $90.4$ $89.8$ $91.7$
COX2 $78.4$ $81.4$ $76.0$ $79.9$ $80.9$ $79.4$ $81.2$ $80.6$ $82.4$
PROTEINS $72.6$ $75.8$ $73.4$ $76.3$ $75.8$ $65.4$ $67.5$ $74.8$ $71.4$ $74.7$ $72.8$ $73.6$
DHFR $78.4$ $81.5$ $70.9$ $81.7$ $80.3$ $82.7$ $80.5$ $81.8$ $83.1$
IMDB-B $72.9$ $71.9$ $73.6$ $73.1$ $74.3$ $54.0$ $68.7$ $71.2$ $74.8$ $69.4$ $68.4$ $69.8$
IMDB-M $50.3$ $47.7$ $52.4$ $50.3$ $52.1$ $36.3$ $46.9$ $48.8$ $47.8$ $46.7$ $43.7$ $44.4$
In line with Carriere et al. (2020), Royer et al. (2021), and Hacquard (2023), we benchmark our performance against the leading methods for these problems, as well as against traditional topological methods documented in the literature. For the latter, P signifies the superior method between the Persistence Images and Persistence Landscapes, while MP indicates the best between Multiparameter Persistence Images and Landscape. All these persistence-based vectorizations are coupled with a Vanillla XGBoost classifier for the learning task. In contrast, the top performing methods for these graph problems predominantly utilize graph kernel methods specifically designed for graph problems. These include: a scale-variant topological method (SV from Tran et al. (2018)), a graph kernel method based on random walks (RetGK from Zhang et al. (2018)), a graph embedding method using spectral distances (FGSD from Verma and Zhang (2017)), a Capsule Graph Neural Network (GCNN from Zhang and Chen (2019)), and a Graph Isomorphism Network (GIN from Xu et al. (2019)). Additionally, we also compare our results with those of PersLay (from Carriere et al. (2020)), a neural network-driven topological vectorization method encapsulating most topological frameworks in the literature, and ATOL (from Royer et al. (2021)), which introduces a fast, learned, unsupervised vectorization method for measures in Euclidean spaces. Lastly, we compare with BBA from Hacquard (2023), which employs rectangle-based classification with boosting on persistence measures. Notably, both PersLay and ATOL use the same eight diagrams with HKS parameters as ours. However, in a few cases for PersLay, the authors employed two HKS diffusion times and then discarded one of them, without any loss in performance.
@@
In Table 1, our results are at the state-of-the-art for the Biomedical benchmark datasets. Our methods consistently surpass all traditional TDA methods, such as P and MP and moreover prove parallel or sometimes better performance in some task with more advanced technique such as ATOL, PersLay, SV and BBA.
@@
Although our methods demonstrate exceptional performance overall, it should be acknowledged that for the Social Networking datasets, they are slightly below the current best methods.

### 05 Jan 2024 v2

# LITE: A Stable Framework for Lattice-Integrated Embedding of Topological Descriptors

@@ Abstract
In this paper, we introduce a new family of descriptors for persistence diagrams. Our approach transforms these diagrams into elements of a finite-dimensional vector space using functionals based on the discrete measures they induce. While our focus is primarily on identity and frequency-based transformations, we do not restrict our approach exclusively to this types of techniques. We term this family of transformations as LITE (Lattice Integrated Topological Embedding) and prove stability for some members of this family against the 1-Kantorovitch-Rubinstein metric, ensuring its responsiveness to subtle data variations. Extensive comparative analysis reveals that our descriptor performs competitively with the current state-of-art from the topological data analysis literature, and often surpasses, the existing methods. This research not only introduces a innovative perspective for data scientists but also critiques the current trajectory of literature on methodologies for vectorizing diagrams. It establishes a foundation for future progress in applying persistence diagrams to data analysis and machine learning under a more simple and effective lens.
Keywords Topological Data Analysis  $\cdot$ Persistent Homology  $\cdot$ Frequency Transformations
@@ 1 Introduction
Topological Data Analysis (TDA) has emerged as a transformative approach in data science, providing useful insights into the underlying structure of complex datasets through the capture of their topological features. The effectiveness of machine learning algorithms, particularly in pattern recognition and feature extraction, underscores the importance of understanding data geometry. TDA offers a more sophisticated exploration of this geometric landscape, leading to numerous successful applications across various fields. Notable examples include neuroscience, where TDA has also been instrumental in understanding complex neural networks and brain connectivity patterns (Anderson et al. (2018), Dabaghian et al. (2012)). In the field of environmental science, TDA has been applied for climate data analysis, as seen in Yamasaki et al. (2009). Additionally, TDA has made significant contributions to materials science, particularly in the study of the structure of materials, as exemplified by Donato et al. (2016). These examples underscore the broad applicability and potential of TDA in enhancing our understanding of complex data across diverse scientific domains.
@@
Our contribution in this study is multi-faceted in the field of TDA. We introduce a novel family of descriptors, termed LITE, which represent a paradigm shift in how persistence diagrams are utilized. Following Divol and Lacombe (2019), we conceptualize persistence diagrams as discrete measures in the +$\mathbb{R}^{2}_{+}$ space, and then transform them into a finite-dimensional vector. This is achieved through the application of functionals to the discrete measures induced by these diagrams. In this work we focus on the identity and frequency based transformations. However, our methodology is not limited to these techniques, allowing for a broader scope of analysis and application. A key aspect of our contribution is the proof of stability of some member of the LITE family against the 1-Kantorovitch-Rubinstein metric. This crucial property underscores the robustness of our approach, ensuring that the LITE descriptors are sensitive and responsive to subtle variations in data.
@@
Our approach not only provides a novel method for vectorizing persistence diagrams but also bridges two different realms: TDA and signal processing theory. This fusion opens the doors to an entirely new perspective in data analysis, offering inspiring ways to unravel complex data patterns and extract meaningful insights. To the best of our knowledge, this is the first instance where tools from classical signal processing theory have been employed to achieve a vectorization method for persistence diagrams. Our method requires minimal tuning and is designed to be user-friendly, making it accessible even to those with limited knowledge of TDA. This novel approach not only simplifies the application of TDA but also expands its potential to a wider audience, promising significant advancements in various fields that rely on complex data analysis. In addition, we also demonstrate how even a simple identity transform within our framework can outperform classical vectorization methods from the TDA literature, such as Persistent Images (Adams et al. (2017)) and Persistence Landscapes (Bubenik (2015)), in common graph classification benchmark tasks. Moreover, this simple transform achieves overall performance comparable to current state-of-the-art methods in the TDA literature on these tasks. This prompts a critique of the direction that current TDA literature has taken in embedding diagrams into vector spaces, highlighting the straightforwardness and simplicity of our approach in comparison.
@@
The remainder of this paper is organized as follows: in Section 2, we review related work that connects TDA and ML, focusing particularly on classical methods of embedding persistence diagrams using vectorization or kernel methods. Section 3 delves into the background of persistent homology and persistent measures, offering foundational knowledge crucial for grasping the ideas from this work. In Section 4, we present our methodology along with the stability theorems that underpin it, providing theoretical insights into our approach. Section 5 is dedicated to showcasing our experimental results. Here, we not only present our findings but also compare them with standard methods from the TDA literature, highlighting the advancements our method brings. Finally, Section 6 is devoted to summarizing our key findings and outlining ideas for further work, suggesting potential avenues for future research and development based upon the groundwork established in this paper. The implementation of LITE has been made available to the public as an open-source resource [^1] https://github.com/majkevh/spectral-master.
@@ 3.1 Persistent Homology
A standard approach for illustrating the insights obtained from persistent homology is the use of a persistence diagram. In our study, we adopt the notation for persistence diagrams as delineated in Cao and Monod (2022). The detailed definition of persistence diagrams is provided in Definition 3.1.
@@
A persistence diagram $D$ is a locally finite multiset of points in the half-plane $\Omega=\{(x,y)\in$ $\left.\mathbb{R}^{2}\mid x\leq y\right\}$ together with points on the diagonal $\partial\Omega=\left\{(x,x)\in\mathbb{R}^{2}\right\}$ counted with infinite multiplicity. Points in $\Omega$ are called off-diagonal points. We denote as persistence of a feature, the value $y-x$.
@@
The collection of all persistence diagrams that possess finite $p$-total persistence is represented by $\mathcal{D}_{p}$. As mentioned in Cao and Monod (2022), the metric space $(\mathcal{D}_{p},\mathrm{W}_{p,q})$ occupies a pivotal role in TDA. Particularly noteworthy is the case when $q=\infty$. In this scenario, the metric space $(\mathcal{D}_{p},\mathrm{W}_{p,\infty})$ constitutes a complete and separable metric space—essentially, a Polish space—for any $1\leq p<\infty$. This characteristic endows the space with significant analytical power, as it ensures that statistical and probabilistic concepts such as probability measures, expectations, and variances are well-defined within this context, as elucidated by Mileyko et al. (2011).
@@ 3.2 Persistence Measures
In the realm of persistence measures, the concept of the optimal partial transport distance plays a pivotal role. This distance has been formally introduced by Figalli and Gigli (2010) and is detailed in Definition 3.6. In our study, we use the same notation as in Cao and Monod (2022).
@@
Building on this foundational concept, we can then consider the metric space defined as $\left(\mathcal{M}^{p},\mathrm{OT}_{p}\right)$. Here, $\mathcal{M}^{p}$ is equipped with the optimal partial transport distance $\mathrm{OT}_{p}$, serving as its metric. As established in Divol and Lacombe (2019), the subset $\mathcal{D}^{p}$ within $\mathcal{M}^{p}$ maintains its closed subspace property and inherits the key characteristics of $\mathcal{M}^{p}$. This includes being a Polish space, applicable for all $1\leq p<\infty$ and $1\leq q\leq\infty$. Consequently, the space $(\mathcal{M}_{p},\mathrm{OT}_{p,q})$ emerges as a robust framework for statistical and probabilistic applications. Moreover, within this framework, $\mathrm{OT}_{p,q}$ coincide with $\mathrm{W}_{p,q}$ when restricted to $\mathcal{D}_{p}$. As also mentioned in Cao and Monod (2022), this alignment underscores the role of persistence measures and the optimal partial transport distance as comprehensive extensions to traditional persistence diagrams and the $p$-Wasserstein distance.
@@ 4 Methodology
In this section, we introduce our framework for embedding persistence measures in vector spaces along with the stability results.
As an overview, the process of vectorizing a persistence diagram using LITE, depicted in Algorithm 1, involves two distinct steps: first, we compute the discretized persistence diagram (refer to Section 4.2), transforming the persistence diagram into a format suitable for further processing. Following this, a functional is applied to this discretized representation to achieve an embedding within a vector space. As detailed in Section 4.1, this work primarily focuses on the trivial functional (with the identity transform) and functionals that define valid frequency and time-frequency distribution functions. Notably, these include the Fourier Transform for frequency distribution, and the Wavelet and Gabor Transforms for time-frequency distribution. These transforms will be introduced in the subsequent subsection. It’s important to note, however, that our framework is versatile and not exclusively limited to these transforms.
Algorithm 1 Lattice Integrated Topological Embedding (LITE)
1:$f$: Transform function with hyperparam., Grid size: $\{N,M\}$, Grid finesse $\Delta$, PDs $\triangleright$ Section 4.1
@@ 4.1 Functionals on Persistence Measures
In this paper, we concentrate on functionals that establish valid frequency and time-frequency distribution functions (hereinafter referred to, for simplicity, as frequency transforms). Both families of these transforms are mapped onto the complex hyperplane (frequency domain from now on). Consequently, $\mathcal{F}(\mathcal{M}^{p})$ can be characterized as a vector space. Furthermore, we also employ the identity transform $f(\mathbf{x},\cdot)=\mathbf{x}$, which effectively maps persistence diagrams directly within the vector space induced by the so called discretized persistence diagrams (refer to Section 4.2).
The choice of selecting functionals that map into the frequency domain is motivated by the fact that persistent diagrams summarize topological features across various length scales of a filtration parameter. This parameter can be seen as a timestamp, indicating when topological features emerge or vanish, therefore it seems natural to interpret them as signals.
In this work, we focus on three possible choices of $f$ for the frequency transforms: two time-frequency distributions, namely the Gabor Transform and the Wavelet Transform, and one frequency transform, specifically the Fourier Transform (which can be regarded as a special case of the Gabor Transform).
@@
4.1.1 Grid-Based Approach in Computational Persistence Theory
Given the practical computational requirements of all considered transforms, they align seamlessly with the framework of our work, which is rooted in the computation of discretized persistence diagrams (for details, refer to Section 4.2). In this framework, measures are confined to allocating mass exclusively at points on a grid (for a definition of the grid measure subspace $\Gamma_{p}$, refer to Lemma 4.1).
Since we are working with finite discrete measures on grids, we employ in this work a discretized version of the transforms discussed in the previous section. For a comprehensive literature review on the Discrete Fourier Transform (DFT), Discrete Wavelet Transform (DWT), and Discrete Gabor Transform (DGT), we direct the reader to Oppenheim and Schafer (2009) and Søndergaard (2007). Given a persistence measure $\mu\in\Gamma_{p}$, the three functionals on the transforms depicted in Section 4.1 are now formulated as follows:
@@ 4.2 Discretized Persistence Diagrams
The concept of discretization of persistence diagrams is the foundation of this work and therefore forms a critical aspect of our methodology.
@@ 4.2.2 Mapping Step
By Proposition 4.2, we therefore obtain a persistence measure $\nu^{\star}$ on $\Gamma_{p}$. We refer to $\nu^{\star}$ as the measure induced by the so-called discretized persistence diagram. Considering the structure of discretized persistence diagrams as previously described, it is straightforward to embed them in a vector space by simply flattening the grid into a vector.
@@ 4.3 Stability
In line with this, we have established stability results with respect to the 1-Wasserstein metric for three of the transforms we considered, namely the identity, Fourier and Gabor Transforms. For the Wavelet Transform, there does not always exist a closed-form expression, which makes finding stability results very challenging. Extending these results to the Wavelet Transform remains a challenge for future work.
@@
Let $\mu=\sum_{\mathbf{x}\in\Theta_{\tau}(D)}\delta_{\mathbf{x}}$ and $\nu=\sum_{\mathbf{x}\in\Theta_{\tau}(D^{\prime})}\delta_{\mathbf{x}}$ be two measures on the discretized persistence diagrams of $D,D^{\prime}$, let $f$ be a functional, either identity, Gabor or Fourier (refer to Section 4.1). Then, there exists a finite constant $c$ such that:
@@ 5 Results
In this section, we demonstrate how LITE effectively preserve topological information, proving to be as competitive as state-of-the-art methods from the TDA literature. Our experiments are conducted on two types of classification tasks: graph-based classification and classification of point clouds derived from dynamical systems. In our experiments, we utilized a Vanilla Random Forest classifier from scikit-learn. The choice of a random forest classifier was driven by two primary considerations: firstly, it facilitates a direct comparison with the study conducted by Royer et al. (2021). Secondly, in line with our objective to showcase the overall effectiveness of our methods relative to existing literature, employing this type of classifier helps to circumvent the necessity for elaborate infrastructure or extensive fine-tuning. Such an approach ensures that the results are not excessively influenced by particular design choices.
@@ 5.1 Graph Classification
The highest accuracies achieved with these transforms (LITE) as well as the accuracy for the identity transform (LITE-IdT) are presented in Table 1. Detailed observations for each transform across the different grids can be found in the Appendix B. Performance is assessed through 10-fold evaluations on each dataset, reporting both average and best 10-fold results.
@@
Dataset ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\star}$ ${}^{\star}$ ${}^{\dagger}$ LITE (Our) LITE-IdT (Our)
${}^{\star}$ ${}^{\dagger}$ ${}^{\star}$ ${}^{\dagger}$
MUTAG $88.3$ $79.2$ $86.1$ $89.8$ $88.3$ $90.4$ $89.8$ $91.7$ $89.2$ $90.7$
COX2 $78.4$ $76.0$ $79.9$ $80.9$ $79.4$ $81.2$ $80.6$ $82.4$ $79.4$ $80.4$
PROTEINS $72.6$ $65.4$ $67.5$ $74.8$ $71.4$ $74.7$ $72.8$ $73.6$ $72.2$ $73.2$
DHFR $78.4$ $70.9$ $81.7$ $80.3$ $82.7$ $80.5$ $81.8$ $83.1$ $81.2$ $82.7$
IMDB-B $72.9$ $54.0$ $68.7$ $71.2$ $74.8$ $69.4$ $68.4$ $69.8$ $67.2$ $68.3$
IMDB-M $50.3$ $36.3$ $46.9$ $48.8$ $47.8$ $46.7$ $43.7$ $44.4$ $43.1$ $44.3$
Table 1: Comparative Analysis of Classification Accuracy with topological methods on Benchmark Graph Datasets. Note: The symbol ${}^{\dagger}$ indicates that the performance should be compared with the Max metric, while the symbol $\star$ denotes that the comparison should be made with the Mean metric.
In line with Carriere et al. (2020), Royer et al. (2021), and Hacquard (2023), we benchmark our performance against the leading methods for these problems, as well as against traditional topological methods documented in the literature. While in Table 1 we only report the comparison with the traditional topological methods, a complete comparison of our methods with also the other non-topological leading methods from the literature is reported in Table 4. For the latter, P signifies the superior method between the Persistence Images and Persistence Landscapes, while MP indicates the best between Multiparameter Persistence Images, Landscapes and kernel embedding methods. All these persistence-based vectorizations are coupled with a Vanillla XGBoost classifier for the learning task. In contrast, the top performing methods for these graph problems predominantly utilize graph kernel methods specifically designed for graph problems. These include: a scale-variant topological method (SV from Tran et al. (2018)), a graph kernel method based on random walks (RetGK from Zhang et al. (2018)), a graph embedding method using spectral distances (FGSD from Verma and Zhang (2017)), a Capsule Graph Neural Network (GCNN from Zhang and Chen (2019)), and a Graph Isomorphism Network (GIN from Xu et al. (2019)). Additionally, we also compare our results with those of PersLay (from Carriere et al. (2020)), a neural network-driven topological vectorization method encapsulating most topological frameworks in the literature, and ATOL (from Royer et al. (2021)), which introduces a fast, learned, unsupervised vectorization method for measures in Euclidean spaces. Lastly, we compare with BBA from Hacquard (2023), which employs rectangle-based classification with boosting on persistence measures. Notably, both PersLay and ATOL use the same eight diagrams with HKS parameters as ours. However, in a few cases for PersLay, the authors employed two HKS diffusion times and then discarded one of them, without any loss in performance.
@@
In Table 1, our results with the frequency transforms are at the state-of-the-art for the Biomedical benchmark datasets. Our methods consistently surpass all traditional TDA methods, such as P and MP and moreover prove parallel or sometimes better performance in some task with more advanced technique such as ATOL, PersLay, SV and BBA.
@@
The accuracy achieved by the identity transform is also noteworthy. While it is generally slightly sub-optimal compared to the frequency transforms we used, it surpasses the performance of other complex vectorization methods on many biomedical datasets. Notably, it outperforms standard vectorizations such as P and MP in many biomedical tasks. This should prompt a reflection on the approach taken by current literature for embedding diagrams. By just discretizing the diagrams on a regular square grid, our method outperformed more intricate techniques.
This observation highlights a significant critique regarding the literature’s direction in embedding diagrams. A simpler and more representative method like ours might be the most effective for preserving information from the diagrams, which also yields an embedding that is directly interpretable and meaningful. This approach suggests that instead of developing increasingly complex vectorization methodologies, straightforward methods may yield more efficient results.
Although our methods demonstrate exceptional performance overall, it should be acknowledged that for the Social Networking datasets, they are slightly below the current best methods. [^3] As previously noted, our framework aligns with that in Royer et al. (2021). We successfully replicated the results reported in their original paper for biomedical datasets on our machine. However, when applying their code to social networking tasks, we observed a performance decrease of approximately $4\%$ compared to the results we present in Table 1. Nonetheless, for consistency, we have included the results from the original paper in our report.